### v1

# Sequences with Inequalities

Abstract.
We consider infinite sequences of positive numbers. The connection between log-concavity and the Bessenrodt–Ono inequality had been in the focus of several papers. This has applications in the white noise distribution theory and combinatorics. We improve a recent result of Benfield and Roy and show that for the sequence of partition numbers $\{p(n)\}$ Nicolas’ log-concavity result implies the result of Bessenrodt and Ono towards $p(n)\,p(m)>p(n+m)$.
Key words and phrases:
Inequalities, Partitions, Sequences
2020 Mathematics Subject Classification:
Primary 05A20, 11P84; Secondary 05A17, 11B83
1. Introduction
Newton ([HLP52], page 104) discovered that the coefficients $\{\alpha(n)\}_{n=0}^{d}$ of polynomials with positive coefficients are log-concave if all the roots are real and therefore are unimodal:
(1) $\alpha(n)^{2}\geq\alpha(n-1)\,\,\,\alpha(n+1),\quad 1\leq n\leq d-1.$
Nevertheless, non-real rooted polynomials as the chromatic polynomials related to the four-color conjecture have this propery [Hu12] and play an important role in algebraic geometry and Hodge theory to combinatorics [Ka22]. Infinite sequences seem to be not accessable by these methods, although, for example $p(n)$ the number of partitions [On04] of $n$, have also turned out to be log-concave [Ni78, DP15] for all even positive $n$ and all odd numbers $n>25$ proven by the circle method of Hardy–Ramanujan and an exact formula by Rademacher. It has been discovered that (1) is also related to the hyperbolicity of the Jensen polynomials associated with the reciprocal of the Dedekind eta function (the generating function of the $p(n)$) of degree two, in the context of studying the Riemann hypothesis [GORZ19]. Recently, another property had been discovered by Bessenrodt and Ono [BO16], which is a mixture of additive and multiplicative behaviour:
$p(n)\,p(m)>p(n+m),\,\,\,n,m\geq 2\text{ and }n+m>9.$
This inequality appears at a first glance as a very special property. But, further sequences turned out to satisfy this Bessenrodt–Ono inequality and similiar inequalities appear as conditions for a white noise distribution theory [AKK00, AKK01]. Initiated by the work by Asai, Kubo, and Kuo there is evidence that under certain conditions log-concavity does imply the Bessenrodt–Ono inequality (BO-inequality), see Corollary 3.
Theorem A (Asai, Kubo, Kuo).
Let $\{\alpha(n)\}_{n=0}^{\infty}$ be a sequence of positive numbers normalized with $\alpha(0)=1$. Then
$\big{\{}\alpha(n)^{2}\leq\alpha(n-1)\alpha(n+1),\,n\geq 1\big{\}}\Longrightarrow\big{\{}\alpha(n)\alpha(m)\leq\alpha(n+m),\,n,m\geq 0\big{\}}.$
We also refer to ([GMU24], Theorem 4.4). Since many important sequences, as $p(n)$, are not log-concave for all $n$, Benfield and Roy [BR24] came up with an interesting result, motivated by the sequence of partition numbers.
Theorem B (Benfield, Roy).
Let $\{\alpha(n)\}_{n=0}^{\infty}$ be a sequence of positive real numbers. Let a natural number $N$ exist, such that for all $n>N$ the sequence is log-concave: $\alpha(n)^{2}\geq\alpha(n-1)\alpha(n+1)$. Let there be a $k\geq 0$ such that the single condition
(2) $\left(\alpha(N+k)\right)^{\frac{1}{N+k}}\geq\left(\alpha(N+k+1)\right)^{\frac{1}{N+k+1}}$
be valid. Then $\alpha(m+n)\leq\alpha(m)\,\alpha(n)$ for all $m,n\geq N+k$.
Benfield and Roy applied their result to the sequence $\{p(n)\}_{n=0}^{\infty}$ Since for $N=26$ and $k=0$ the condition (2) is fulfilled, they obtain from the log-concavity property that the BO-inequality holds for $m,n>25$.
Unfortunately, this does not cover the complete result by Bessenrodt–Ono. There are still infinitely many pairs $(m,n)$ left to be checked. For example the BO-inequality holds true for all pairs $(2,n)$, where $n>25$.
In this paper, motivated by the result by Benfield and Roy, we give a new criterion, again, providing a sufficient condition (see Example 1), but strong enough to imply the complete result by Bessenrodt and Ono (see Corollary 2):
Theorem 1.
Let a sequence $\{\alpha(n)\}_{n=0}^{\infty}$ with $\alpha(n)>0$ be given. Suppose there is an $n_{0}\geq 1$ such for all $n\geq n_{0}$ the sequence is log-concave, $\alpha(n)^{2}\geq\alpha(n-1)\,\alpha(n+1)$, for all $n\geq n_{0}$ and satisfies
(3) $\alpha(n_{0})^{1/n_{0}}>\frac{\alpha(n_{0})}{\alpha(n_{0}-1)}.$
Let
$A=\left\{1\leq a\leq n_{0}-2\,:\,\alpha(a)^{1/a}>\frac{\alpha({n_{0}})}{\alpha({n_{0}-1})}\right\}$
and let further $m,n\geq 1$ satisfy
$\big{(}m,n\geq n_{0}-1\big{)}\,\vee\,\big{(}\,m\in A,n\geq n_{0}-1\big{)}\,\vee\,\big{(}m\geq n_{0}-1,n\in A\big{)}.$
Then the Bessenrodt–Ono inequality $\alpha(m)\,\alpha(n)>\alpha(m+n)$ is satisfied.
[…]
Corollary 2.
Let the sequence $\{p(n)\}$ of partition numbers be given. Then the log-concavity result by Nicolas implies the BO-inequality result by Bessenrodt and Ono.
Essentially Theorem 1 is still true if we assume
(4) $\alpha(n_{0})^{1/n_{0}}\geq\frac{\alpha(n_{0})}{\alpha(n_{0}-1)}$
and put
$A=\left\{1\leq a\leq n_{0}-2\,:\,\alpha(a)^{1/a}\geq\frac{\alpha({n_{0}})}{\alpha({n_{0}-1})}\right\}.$
If we suppose $m,n\geq 1$ satisfy
$\big{(}m,n\geq n_{0}-1\big{)}\,\vee\,\big{(}\,m\in A,n\geq n_{0}-1\big{)}\,\vee\,\big{(}m\geq n_{0}-1,n\in A\big{)},$
then
(5) $\alpha(m)\,\alpha(n)\geq\alpha(m+n).$
2. Proof of Theorem 1
For our purposes we quickly derive the following from the proof of Theorem A in [AKK00].
Corollary 3 (to the proof of Theorem A).
Let $\left\{\alpha\left({n}\right)\right\}_{n=0}^{\infty}$ be a sequence of positive numbers with $\alpha\left({0}\right)>1$. If $\left\{\alpha\left({n}\right)\right\}_{n=0}^{\infty}$ is log-concave, then
(6) $\alpha\left({n}\right)\alpha\left({m}\right)>\alpha\left({n+m}\right),\quad\forall n,m\geq 0.$
Proof.
Let $\alpha\left(n\right)$ be the elements of a sequence as in Theorem A. For such a sequence the authors of [AKK00] arrive for $n\geq 0$ and $m\geq 1$ in the first line on page 84 at the inequality $\alpha\left(n\right)\alpha\left(m\right)\leq\alpha\left(0\right)\alpha\left(n+m\right)$. If we now assume $\alpha\left(0\right)<1$ (instead of $\alpha\left(0\right)=1$) this shows
(7) $\alpha\left(n\right)\alpha\left(m\right)<\alpha\left(n+m\right).$
Now we consider a sequence consisting of $\alpha\left({n}\right)$ with assumptions as stated in the corollary. Since our sequence is log-concave if we put $\beta\left(n\right)=1/\alpha\left({n}\right)$ this sequence will be log-convex and we obtain from (7) by inversion $\alpha\left({n}\right)\alpha\left({m}\right)>\alpha\left({n+m}\right)$. ∎
With this auxiliary result we can now finish the proof of Theorem 1.
Proof of Theorem 1.
Firstly, let $a,b\geq n_{0}-1$. Then we define
$\beta\left({n}\right)=\left\{\begin{array}[]{ll}\left(\frac{\alpha\left({n_{0}-1}\right)}{\alpha\left({n_{0}}\right)}\right)^{n_{0}-n}\alpha\left({n_{0}}\right),&n<n_{0},\\ \alpha\left({n}\right),&n\geq n_{0}.\end{array}\right.$
Then $\beta\left({0}\right)=\left(\frac{\alpha\left({n_{0}-1}\right)}{\alpha\left({n_{0}}\right)}\right)^{n_{0}}\alpha\left({n_{0}}\right)>1$ by (3). The complete sequence consisting of the $\beta\left({n}\right)$ is now log-concave and we can apply the corollary of the result by Asai, Kubo, and Kuo [AKK00]. From Corollary 3 with $\beta\left({0}\right)>1$ we obtain $\beta\left({a}\right)\beta\left({b}\right)>\beta\left({a+b}\right)$ for all $a,b\geq 0$. Since $\beta\left({n}\right)=\alpha\left({n}\right)$ for all $n\geq n_{0}-1$ the Bessenrodt–Ono inequality (6) is also fulfilled for all $a,b\geq n_{0}-1$.
We have $\left(\alpha\left({n}\right)\right)^{2}\geq\alpha\left({n-1}\right)\alpha\left({n+1}\right)$ for $n\geq n_{0}$. Since all $\alpha\left({n}\right)$ are positive we have
$\frac{\alpha\left({n}\right)}{\alpha\left({n-1}\right)}\geq\frac{\alpha\left({n+1}\right)}{\alpha\left({n}\right)}.$
Now we assume $a\in A$ and $b\geq n_{0}-1$. Then
$\frac{\alpha\left({a+b}\right)}{\alpha\left({b}\right)}=\prod_{k=b+1}^{a+b}\frac{\alpha\left({k}\right)}{\alpha\left({k-1}\right)}\leq\left(\frac{\alpha\left({n_{0}}\right)}{\alpha\left({n_{0}-1}\right)}\right)^{a}.$
Therefore, by our assumption $\frac{\alpha\left({a+b}\right)}{\alpha\left({b}\right)}<\alpha\left({a}\right)$ which implies $\alpha\left({a+b}\right)<\alpha\left({a}\right)\alpha\left({b}\right)$. ∎
3. Applications and variants
The following first example shows that our conditions $\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ and $\left(\alpha\left(a\right)\right)^{1/a}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ is sufficient but not necessary.
Example 1.
Let $n_{0}=4$, $\alpha\left({0}\right)$ and $\alpha\left({1}\right)$ positive but arbitrary, $\alpha\left({2}\right)=7$, $\alpha\left({3}\right)=5$, $\alpha\left({4}\right)=15$, and $\alpha\left({n}\right)=\frac{30}{\left(n-4\right)!}$ for $n\geq 5$. We have $\frac{\left(\alpha\left({3}\right)\right)^{2}}{\alpha\left({2}\right)\alpha\left({4}\right)}=\frac{25}{105}<1$, $\frac{\left(\alpha\left({4}\right)\right)^{2}}{\alpha\left({3}\right)\alpha\left({5}\right)}=\frac{225}{150}>1$, $\frac{\left(\alpha\left(5\right)\right)^{2}}{\alpha\left(4\right)\alpha\left(6\right)}=\frac{900}{225}>1$, and $\frac{\left(\alpha\left({n}\right)\right)^{2}}{\alpha\left({n-1}\right)\alpha\left({n+1}\right)}=\frac{n-3}{n-4}>1$ for $n\geq 6$. The sequence consisting of the $\alpha\left({n}\right)$ is log-concave for $n\geq n_{0}=4$. Furthermore, $\left(\alpha\left(2\right)\right)^{2}=49>15=\alpha\left(4\right)$, $\alpha\left({2}\right)\alpha\left({3}\right)=35>30=\alpha\left({5}\right)$, $\alpha\left({2}\right)\alpha\left({4}\right)=105>15=\alpha\left({6}\right)$, $\alpha\left({2}\right)\alpha\left({b}\right)=\frac{210}{\left(b-4\right)!}>\frac{30}{\left(b-2\right)!}=\alpha\left({2+b}\right)$ for $b\geq 5$, $\left(\alpha\left(3\right)\right)^{2}=25>15=\alpha\left(6\right)$, $\alpha\left(3\right)\alpha\left(4\right)=75>5=\alpha\left(7\right)$, $\alpha\left(3\right)\alpha\left(b\right)=\frac{150}{\left(b-4\right)!}>\frac{30}{\left(b-1\right)!}=\alpha\left(3+b\right)$ for $b\geq 5$, $\left(\alpha\left(4\right)\right)^{2}=225>\frac{5}{4}=\alpha\left(8\right)$, $\alpha\left(4\right)\alpha\left(b\right)=\frac{450}{\left(b-4\right)!}>\frac{30}{b!}=\alpha\left(4+b\right)$ for $b\geq 5$, and $\alpha\left(a\right)\alpha\left(b\right)=\frac{900}{\left(a-4\right)!\left(b-4\right)!}>\frac{30}{\left(a+b-4\right)!}=\alpha\left(a+b\right)$ for $a,b\geq 5$. But $\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}=15^{1/4}<3=\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ and $\left(\alpha\left({2}\right)\right)^{1/2}=\sqrt{7}<3=\frac{\alpha\left({n_{0}}\right)}{\alpha\left({n_{0}-1}\right)}$.
On the other hand we have the following in the opposite direction.
Proposition 4.
Let $\alpha\left({n}\right)$ be the positive elements of a sequence that is log-concave for all $n\geq 1$ and which satisfies the Bessenrodt–Ono inequality (5)
$\alpha\left(n\right)\alpha\left(m\right)\geq\alpha\left(n+m\right)$
for all $n,m\geq 0$. Then $\left(\alpha\left({n}\right)\right)^{1/n}\geq\frac{\alpha\left({n}\right)}{\alpha\left({n-1}\right)}$ for all $n\geq 1$.
Proof.
We have $\alpha\left({n}\right)=\alpha\left({0}\right)\prod_{k=1}^{n}\frac{\alpha\left({k}\right)}{\alpha\left({k-1}\right)}\geq\left(\frac{\alpha\left({n}\right)}{\alpha\left({n-1}\right)}\right)^{n}$ as $\frac{\alpha\left({k}\right)}{\alpha\left({k-1}\right)}\geq\frac{\alpha\left({k+1}\right)}{\alpha\left({k}\right)}$. ∎
Example 2.
Now we consider the plane partition numbers $\mathop{\rm pp}\left(n\right)$. By [OPR22] this sequence is log-concave for $n\geq n_{0}=12$. We can check that $\left(\mathop{\rm pp}\left(12\right)\right)^{1/12}>\frac{1479}{859}$. We obtain $\left\{1\leq a\leq 10:\left(\mathop{\rm pp}\left(a\right)\right)^{1/a}>\frac{\mathop{\rm pp}\left(12\right)}{\mathop{\rm pp}\left(11\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 10\right\}$. Computationally checking the cases for $2\leq a,b\leq 10$ then yields our previous result with Tröger, see [HNT23].
Example 3.
Let $\sum_{n=0}^{\infty}\bar{p}\left(n\right)q^{n}=\prod_{n=1}^{\infty}\frac{1+q^{n}}{1-q^{n}}$ define the overpartitions. By the result of Engel [En17] it is strictly log-concave $\left(\bar{p}\left(n\right)\right)^{2}>\bar{p}\left(n-1\right)\bar{p}\left(n+1\right)$ for $n\geq 3$. Nevertheless it is already log-concave (1) for all $n\geq 1$. We use $n_{0}=4$. Then $\left(\bar{p}\left(n_{0}\right)\right)^{1/n_{0}}>\frac{7}{4}$ and $A=\left\{1\leq a\leq 2:\left(\bar{p}\left(a\right)\right)^{1/a}>\frac{\bar{p}\left(4\right)}{\bar{p}\left(3\right)}\right\}=\left\{1,2\right\}$. By our result we obtain that the BO-inequality (6) is fulfilled for all $n\geq 3$ or $m\geq 3$. By computationally checking the cases $1\leq n,m\leq 2$ we obtain the result of [LZ21]: the overpartitions fulfill the BO-inequality (6) for all $n,m\geq 1$ except for $\left(n,m\right)\in\left\{\left(1,1\right),\left(1,2\right),\left(2,1\right)\right\}$ where we have equality. Equality also holds when $n=0$ or $m=0$.
Example 4.
Let $\sum_{n=0}^{\infty}p_{k}\left(n\right)q^{n}=\frac{\prod_{n=1}^{\infty}\left(1-q^{kn}\right)}{\prod_{n=1}^{\infty}\left(1-q^{n}\right)}$ define the $k$-regular partitions.
  1. (1)
    Let $k=2$. By a result by Dong and Ji [DJ24] this is log-concave for $n\geq n_{0}=33$. We obtain $\left(p_{2}\left(33\right)\right)^{1/33}>\frac{224}{195}$ and $A=\left\{1\leq a\leq 31:\left(p_{2}\left(a\right)\right)^{1/a}>\frac{p_{2}\left(33\right)}{p_{2}\left(32\right)}\right\}=\left\{a\in\mathbb{N}_{0}:3\leq a\leq 31\right\}$. By computationally checking the cases $3\leq n,m\leq 31$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the BO-inequality.
  2. (2)
    Now let $k=3$. Then again by [DJ24] this is log-concave for $n\geq n_{0}=58$. We obtain $\left(p_{3}\left(58\right)\right)^{1/58}>\frac{525}{463}$ and $A=\left\{1\leq a\leq 56:\left(p_{3}\left(a\right)\right)^{1/a}>\frac{p_{3}\left(58\right)}{p_{3}\left(57\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 56\right\}$. Computationally checking the cases $2\leq n,m\leq 56$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the BO-inequality.