### v2

# Sequences with Inequalities

Abstract.
We consider infinite sequences of positive numbers. The connection between log-concavity and the Bessenrodt–Ono inequality had been in the focus of several papers. This has applications in the white noise distribution theory and combinatorics. We improve a recent result by Benfield and Roy and show that for the sequence of partition numbers $\{p(n)\}$ Nicolas’ log-concavity result implies the result by Bessenrodt and Ono towards p(n)p(m)>p(n+m)p(n)\,p(m)>p(n+m). We provide several examples. Benfield and Roy gave a conjecture related to $\ell$-ary partition numbers. We prove part of this conjecture.
Key words and phrases:
Inequalities, Partitions, Sequences
2020 Mathematics Subject Classification:
Primary 05A20, 11P84; Secondary 05A17, 11B83
Corresponding author
1. Introduction
Newton ([HLP52], page 104) discovered that the coefficients $\{\alpha(n)\}_{n=0}^{d}$ of polynomials with positive coefficients are log-concave if all the roots are real and therefore are unimodal:
(1) $\alpha(n)^{2}\geq\alpha(n-1)\,\,\,\alpha(n+1),\quad 1\leq n\leq d-1.$
Nevertheless, non-real rooted polynomials as the chromatic polynomials related to the four-color conjecture have this propery [Hu12] and play an important role in algebraic geometry and Hodge theory to combinatorics [Ka22]. Infinite sequences seem to be not accessable by these methods, although, for example $p(n)$ the number of partitions [On04] of $n$, have also turned out to be log-concave [Ni78, DP15] for all even positive $n$ and all odd numbers $n>25$ proven by the circle method by Hardy–Ramanujan and an exact formula by Rademacher. It has been discovered that (1) is also related to the hyperbolicity of the Jensen polynomials associated with the reciprocal of the Dedekind eta function (the generating function of the $p(n)$) of degree two, in the context of studying the Riemann hypothesis [GORZ19]. Recently, another property had been discovered by Bessenrodt and Ono [BO16], which is a mixture of additive and multiplicative behaviour:
$p(n)\,p(m)>p(n+m),\,\,\,n,m\geq 2\text{ and }n+m>9.$
This inequality appears at a first glance as a very special property. However, additional sequences were found to satisfy this Bessenrodt–Ono inequality and similiar inequalities appear as conditions for a white noise distribution theory [AKK00, AKK01]. Initiated by the work by Asai, Kubo, and Kuo there is evidence that under certain conditions log-concavity does imply the Bessenrodt–Ono inequality, see Corollary 6.
Theorem A (Asai, Kubo, Kuo).
Let $\{\alpha(n)\}_{n=0}^{\infty}$ be a sequence of positive numbers normalized with $\alpha(0)=1$. Then
$\big\{\alpha(n)^{2}\leq\alpha(n-1)\alpha(n+1),\,n\geq 1\big\}\Longrightarrow\big\{\alpha(n)\alpha(m)\leq\alpha(n+m),\,n,m\geq 0\big\}.$
We also refer to ([GMU25], Theorem 4.4). Since many important sequences, as $p(n)$, are not log-concave for all $n$, Benfield and Roy [BR24a] came up with an interesting result, motivated by the sequence of partition numbers.
Theorem B (Benfield, Roy).
Let $\{\alpha(n)\}_{n=0}^{\infty}$ be a sequence of positive real numbers. Let a natural number $N$ exist, such that for all $n>N$ the sequence is log-concave: $\alpha(n)^{2}\geq\alpha(n-1)\alpha(n+1)$. Let there be a $k\geq 0$ such that the single condition
(2) $\left(\alpha(N+k)\right)^{\frac{1}{N+k}}\geq\left(\alpha(N+k+1)\right)^{\frac{1}{N+k+1}}$
be valid. Then $\alpha(m+n)\leq\alpha(m)\,\alpha(n)$ for all $m,n\geq N+k$.
Benfield and Roy applied their result to the sequence $\{p(n)\}_{n=0}^{\infty}$. Since for $N=26$ and $k=0$ the condition (2) is fulfilled, they obtain from the log-concavity property that the Bessenrodt–Ono inequality holds for $m,n>25$.
Unfortunately, this did not cover the complete result by Bessenrodt–Ono. There are still infinitely many pairs $(m,n)$ left to be checked. For example the Bessenrodt–Ono inequality holds true for all pairs $(2,n)$, where $n>25$.
In this paper, motivated by the result by Benfield and Roy, we give a new criterion, again, providing a sufficient condition (see Example 7), but strong enough to imply the complete result by Bessenrodt and Ono (see Corollary 5):
Theorem 1.
Let a sequence $\{\alpha(n)\}_{n=0}^{\infty}$ with $\alpha(n)>0$ be given. Suppose there is an $n_{0}\geq 1$ such for all $n\geq n_{0}$ the sequence is log-concave, $\alpha(n)^{2}\geq\alpha(n-1)\,\alpha(n+1)$, for all $n\geq n_{0}$ and satisfies
(3) $\alpha(n_{0})^{1/n_{0}}>\frac{\alpha(n_{0})}{\alpha(n_{0}-1)}.$
Let
$A=\left\{1\leq a\leq n_{0}-2\,:\,\alpha(a)^{1/a}>\frac{\alpha({n_{0}})}{\alpha({n_{0}-1})}\right\}$
and let further $m,n\geq 1$ satisfy
$\big(m,n\geq n_{0}-1\big)\,\vee\,\big(\,m\in A,n\geq n_{0}-1\big)\,\vee\,\big(m\geq n_{0}-1,n\in A\big).$
Then the Bessenrodt–Ono inequality $\alpha(m)\,\alpha(n)>\alpha(m+n)$ is satisfied.
Remark 2.
  1. (1)
    The index $n_{0}$ is not unique and may be chosen suitably.
  2. (2)
    After Theorem 1 was published on arXiv, see [HN24], an improved version [BR24b] of Benfield and Roy’s result appeared also covered by our result.
We illustrate the previous theorem with two examples.
Example 3 (see ([BR24b], 1.1, ex. 4)).
Let $\{\alpha(n)\}$ be given by $\alpha\left(0\right)=1/2$ and $\alpha\left(n\right)=2^{\sqrt{n}}$. Then the sequence is log-concave for all $n\geq 1$. Let $n_{0}=2$. Then $\alpha(2)^{1/2}>\alpha(2)/\alpha(1)$. Therefore, condition (3) is fulfilled and we obtain the strict Bessenrodt–Ono inequality as stated in Theorem 1 for $n,m\geq n_{0}-1=1$. In this case, the set AA is empty and plays no role.
Example 4 (see [BR24b], 1.1, ex. 5)).
Define the sequence $\alpha\left(n\right)=\left(1.24\right)^{n}+0.000001$ for $0\leq n\leq 24$ and $\alpha\left(n\right)=p\left(n\right)$ for $n\geq 25$. This sequence is log-concave for $n\geq 26$. Let $n_{0}=27$. Then condition (3) is fulfilled and $A=\left\{2,3,\ldots,25\right\}$. This leads to the strict Bessenrodt–Ono inequality for all $n,m\geq 2$ with $n\geq 26$ or $m\geq n_{0}-1=26$.
Note that taking $n_{0}=26$ we obtain the strict Bessenrodt–Ono inequality only for all $n,m\geq 25$.
In the case of partition numbers $p(n)$ Theorem 1 leads to the following interesting application.
Corollary 5.
Let the sequence $\{p(n)\}$ of partition numbers be given. Then the log-concavity result by Nicolas [Ni78] implies the Bessenrodt–Ono inequality result by Bessenrodt and Ono [BO16].
Proof.
Let $\alpha(n)=p(n)$ and $n_{0}=26$. Then (3) is satisfied. Moreover,
$A=\{a\in\mathbb{N}_{0}:2\leq a<25\}.$
Therefore, only finitely many pairs given by $\{\left(m,n\right)\in\mathbb{N}_{0}^{2}\,:\,2\leq m,n<25\}$ remain to be checked by a computer calculation. ∎
The sequence $\{c^{n}\}$ for $c>0$ is log-concave but does not satisfy the strict Bessenrodt–Ono inequality. Nevertheless, a minor modification of Theorem 1 can be applied. Let
$\alpha(n_{0})^{1/n_{0}}\geq\frac{\alpha(n_{0})}{\alpha(n_{0}-1)}$
and
$A:=\left\{1\leq a\leq n_{0}-2\,:\,\alpha(a)^{1/a}\geq\frac{\alpha({n_{0}})}{\alpha({n_{0}-1})}\right\}.$
If $m,n\geq 1$ satisfy
$\big(m,n\geq n_{0}-1\big)\,\vee\,\big(\,m\in A,n\geq n_{0}-1\big)\,\vee\,\big(m\geq n_{0}-1,n\in A\big),$
then
(4) $\alpha(m)\,\alpha(n)\geq\alpha(m+n).$
In the last section we invest into a recent conjecture by Benfield and Roy related to $\ell$-ary partition numbers [BR24b].
2. Proof of Theorem 1
For our purposes we quickly derive the following from the proof of Theorem A in [AKK00].
Corollary 6 (to the proof of Theorem A).
Let $\left\{\alpha\left({n}\right)\right\}_{n=0}^{\infty}$ be a sequence of positive numbers with $\alpha\left({0}\right)>1$. If $\left\{\alpha\left({n}\right)\right\}_{n=0}^{\infty}$ is log-concave, then
(5) $\alpha\left({n}\right)\alpha\left({m}\right)>\alpha\left({n+m}\right),\quad\forall n,m\geq 0.$
Proof.
Let $\alpha\left(n\right)$ be the members of a sequence as in Theorem A. For such a sequence the authors of [AKK00] arrive for $n\geq 0$ and $m\geq 1$ in the first line on page 84 at the inequality $\alpha\left(n\right)\alpha\left(m\right)\leq\alpha\left(0\right)\alpha\left(n+m\right)$. If we now assume $\alpha\left(0\right)<1$ (instead of $\alpha\left(0\right)=1$) this shows
(6) $\alpha\left(n\right)\alpha\left(m\right)<\alpha\left(n+m\right).$
Now we consider a sequence consisting of $\alpha\left({n}\right)$ with assumptions as stated in the corollary. Since our sequence is log-concave if we put $\beta\left(n\right)=1/\alpha\left({n}\right)$ this sequence will be log-convex and we obtain from (6) by inversion $\alpha\left({n}\right)\alpha\left({m}\right)>\alpha\left({n+m}\right)$. ∎
With this auxiliary result we can now finish the proof of Theorem 1.
Proof of Theorem 1.
Firstly, let $a,b\geq n_{0}-1$. Then we define
$\beta\left({n}\right)=\left\{\begin{array}[]{ll}\left(\frac{\alpha\left({n_{0}-1}\right)}{\alpha\left({n_{0}}\right)}\right)^{n_{0}-n}\alpha\left({n_{0}}\right),&n<n_{0},\\ \alpha\left({n}\right),&n\geq n_{0}.\end{array}\right$.
Then $\beta\left({0}\right)=\left(\frac{\alpha\left({n_{0}-1}\right)}{\alpha\left({n_{0}}\right)}\right)^{n_{0}}\alpha\left({n_{0}}\right)>1$ by (3). The complete sequence consisting of the $\beta\left({n}\right)$ is now log-concave and we can apply the corollary of the result by Asai, Kubo, and Kuo [AKK00]. From Corollary 6 with $\beta\left({0}\right)>1$ we obtain $\beta\left({a}\right)\beta\left({b}\right)>\beta\left({a+b}\right)$ for all $a,b\geq 0$. Since $\beta\left({n}\right)=\alpha\left({n}\right)$ for all $n\geq n_{0}-1$ the Bessenrodt–Ono inequality (5) is also fulfilled for all $a,b\geq n_{0}-1$.
We have $\left(\alpha\left({n}\right)\right)^{2}\geq\alpha\left({n-1}\right)\alpha\left({n+1}\right)$ for $n\geq n_{0}$. Since all $\alpha\left({n}\right)$ are positive we have
$\frac{\alpha\left({n}\right)}{\alpha\left({n-1}\right)}\geq\frac{\alpha\left({n+1}\right)}{\alpha\left({n}\right)}.$
Now we assume $a\in A$ and $b\geq n_{0}-1$. Then
$\frac{\alpha\left({a+b}\right)}{\alpha\left({b}\right)}=\prod_{k=b+1}^{a+b}\frac{\alpha\left({k}\right)}{\alpha\left({k-1}\right)}\leq\left(\frac{\alpha\left({n_{0}}\right)}{\alpha\left({n_{0}-1}\right)}\right)^{a}.$
Therefore, by our assumption $\frac{\alpha\left({a+b}\right)}{\alpha\left({b}\right)}<\alpha\left({a}\right)$ which implies $\alpha\left({a+b}\right)<\alpha\left({a}\right)\alpha\left({b}\right)$. ∎
3. Applications and variants
The following example shows that our conditions
$\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}\text{ and }\left(\alpha\left(a\right)\right)^{1/a}>\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$
are sufficient but not necessary to obtain the Bessenrodt–Ono inequality.
Example 7.
We consider the sequence given in Table 1 with $\alpha\left({0}\right)$ and $\alpha\left({1}\right)$ positive.
[…]
Table 1. Definition of the sequence from Example 7
The sequence consisting of the $\alpha\left({n}\right)$ is log-concave for $n\geq n_{0}=4$, see Table 2.
[…]
Table 2. Log-concavity and -convexity in Example 7
Furthermore, we obtain the Bessenrodt–Ono inequality for all instances $a,b\geq 2$, see Table 3.
[…]
Table 3. Instances of the Bessenrodt–Ono inequality $\alpha\left(a\right)\alpha\left(b\right)>\alpha\left(a+b\right)$ for Example 7
But for $n_{0}=4$ we obtain $\left(\alpha\left(n_{0}\right)\right)^{1/n_{0}}=15^{1/4}<3=\frac{\alpha\left(n_{0}\right)}{\alpha\left(n_{0}-1\right)}$ and $\left(\alpha\left({2}\right)\right)^{1/2}=\sqrt{7}<3=\frac{\alpha\left({n_{0}}\right)}{\alpha\left({n_{0}-1}\right)}$.
On the other hand we have the following in the opposite direction.
Proposition 8.
Let $\alpha\left({n}\right)$ be a sequence of positive numbers that is log-concave for all $n\geq 1$ and which satisfies the Bessenrodt–Ono inequality (4)
$\alpha\left(n\right)\alpha\left(m\right)\geq\alpha\left(n+m\right)$
for all $n,m\geq 0$. Then $\left(\alpha\left({n}\right)\right)^{1/n}\geq\frac{\alpha\left({n}\right)}{\alpha\left({n-1}\right)}$ for all $n\geq 1$.
Proof.
We have $\alpha\left({n}\right)=\alpha\left({0}\right)\prod_{k=1}^{n}\frac{\alpha\left({k}\right)}{\alpha\left({k-1}\right)}\geq\left(\frac{\alpha\left({n}\right)}{\alpha\left({n-1}\right)}\right)^{n}$ as $\frac{\alpha\left({k}\right)}{\alpha\left({k-1}\right)}\geq\frac{\alpha\left({k+1}\right)}{\alpha\left({k}\right)}$. ∎
Example 9.
Now we consider the plane partition numbers $\mathop{\rm pp}\left(n\right)$. By [OPR22] this sequence is log-concave for $n\geq n_{0}=12$. We can check that $\left(\mathop{\rm pp}\left(12\right)\right)^{1/12}>\frac{\mathop{\rm pp}\left(12\right)}{\mathop{\rm pp}\left(11\right)}$. We obtain
$A=\left\{1\leq a\leq 10:\left(\mathop{\rm pp}\left(a\right)\right)^{1/a}>\frac{\mathop{\rm pp}\left(12\right)}{\mathop{\rm pp}\left(11\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 10\right\}.$
Computationally checking the cases for $2\leq m,n\leq 10$ then yields our previous result with Tröger, see [HNT23].
Example 10.
Let $\sum_{n=0}^{\infty}\bar{p}\left(n\right)q^{n}=\prod_{n=1}^{\infty}\frac{1+q^{n}}{1-q^{n}}$ define the overpartitions. By the result by Engel [En17] it is strictly log-concave $\left(\bar{p}\left(n\right)\right)^{2}>\bar{p}\left(n-1\right)\bar{p}\left(n+1\right)$ for $n\geq 3$. Nevertheless it is already log-concave (1) for all $n\geq 1$. Let $n_{0}=4$. Then $\left(\bar{p}\left(n_{0}\right)\right)^{1/n_{0}}>\frac{\bar{p}\left(4\right)}{\bar{p}\left(3\right)}$ and
$A=\left\{1\leq a\leq 2:\left(\bar{p}\left(a\right)\right)^{1/a}>\frac{\bar{p}\left(4\right)}{\bar{p}\left(3\right)}\right\}=\left\{1,2\right\}.$
By our result we obtain that the Bessenrodt–Ono inequality (5) is fulfilled for all $n\geq 3$ or $m\geq 3$. By computationally checking the cases $1\leq n,m\leq 2$ we obtain the result of [LZ21]: the overpartitions fulfill the Bessenrodt–Ono inequality (5) for all $n,m\geq 1$ except for $\left(n,m\right)\in\left\{\left(1,1\right),\left(1,2\right),\left(2,1\right)\right\}$ where we have equality. Equality also holds when $n=0$ or $m=0$.
Example 11.
Let $\sum_{n=0}^{\infty}p_{k}\left(n\right)q^{n}=\frac{\prod_{n=1}^{\infty}\left(1-q^{kn}\right)}{\prod_{n=1}^{\infty}\left(1-q^{n}\right)}$ define the $k$-regular partitions.
  1. (1)
    Let $k=2$. By a result by Dong and Ji [DJ24] this is log-concave for $n\geq n_{0}=33$. We obtain $\left(p_{2}\left(33\right)\right)^{1/33}>\frac{p_{2}\left(33\right)}{p_{2}\left(32\right)}$ and
    $A=\left\{1\leq a\leq 31:\left(p_{2}\left(a\right)\right)^{1/a}>\frac{p_{2}\left(33\right)}{p_{2}\left(32\right)}\right\}=\left\{a\in\mathbb{N}_{0}:3\leq a\leq 31\right\}.$
    By computationally checking the cases $3\leq n,m\leq 31$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the Bessenrodt–Ono inequality.
  2. (2)
    Now let $k=3$. Then again by [DJ24] this is log-concave for $n\geq n_{0}=58$. We obtain $\left(p_{3}\left(58\right)\right)^{1/58}>\frac{p_{3}\left(58\right)}{p_{3}\left(57\right)}$ and
    $A=\left\{1\leq a\leq 56:\left(p_{3}\left(a\right)\right)^{1/a}>\frac{p_{3}\left(58\right)}{p_{3}\left(57\right)}\right\}=\left\{a\in\mathbb{N}_{0}:2\leq a\leq 56\right\}.$
    Computationally checking the cases $2\leq n,m\leq 56$ we obtain the result by Beckwith and Bessenrodt [BB16] regarding the Bessenrodt–Ono inequality.
Proposition 12.
Let $\left\{\alpha\left(n\right)\right\}_{n=0}^{\infty}$ be a sequence of positive numbers such that
$\left(\alpha\left(n\right)\right)^{\frac{1}{n}}>\left(\alpha\left(n+1\right)\right)^{\frac{1}{n+1}}$
for all $n\geq n_{0}\geq 1$. Then $\alpha\left(n\right)\alpha\left(m\right)>\alpha\left(n+m\right)$ for all $n,m\geq n_{0}$.
Proof.
We obtain via a direct computation
$\alpha\left(n\right)\alpha\left(m\right)$ $=$ $\left(\left(\alpha\left(n\right)\right)^{\frac{1}{n}}\right)^{n}\left(\left(\alpha\left(m\right)\right)^{\frac{1}{m}}\right)^{m}$
$>$ $\left(\left(\alpha\left(n+m\right)\right)^{\frac{1}{n+m}}\right)^{n}\left(\left(\alpha\left(n+m\right)\right)^{\frac{1}{n+m}}\right)^{m}=\alpha\left(n+m\right).$
4. On a conjecture by Benfield and Roy
Benfield and Roy made a conjecture ([BR24b], Conjecture 5.1) on the strict Bessenrodt–Ono inequality of $\ell$-ary partition numbers $b^{\ell}\left(n\right)$ of $n$. It is well known ([GMU25], Example 4.7) that the $\ell$-ary sequences are not log-concave. We first invest in the case $\ell=5$. Obviously $b^{5}\left(n\right)b^{5}\left(m\right)\leq b^{5}\left(n+m\right)$ if $m$ or $n\in\left\{1,2,3,4\right\}$. This is similar to case $m$ or $n\in\left\{1\right\}$ in [BO16]. For the other cases they conjectured that $7\leq m\leq 9$ and $6\leq n\leq 9$. But since $b^{5}\left(6\right)=2=b^{5}\left(7\right)$ and $b^{5}\left(13\right)=3$ this has to be modified. The modified conjecture is recorded in Table 4.
[…]
Table 4. Conjectured equality and failure of $b^{5}\left(n\right)b^{5}\left(m\right)\geq b^{5}\left(n+m\right)$ for $n\geq m\geq 1$
Theorem 13.
The conjecture holds true for all pairs $m,n\leq 7500$.
This has been verified with PARI/GP.
The same applies to the $\ell$-ary partition numbers $b^{\ell}\left(n\right)$ (Table 5).
$b^{\ell}\left(n\right)b^{\ell}\left(m\right)=b^{\ell}\left(n+m\right)$ $b^{\ell}\left(n\right)b^{\ell}\left(m\right)<b^{\ell}\left(n+m\right)$
$1\leq m\leq\ell-1$ with $1\leq m\leq\ell-1$ with
$n\equiv 0,1,\ldots,\ell-m-1\pmod{\ell}$ $n\equiv\ell-m,\ldots,\ell-1\pmod{\ell}$
$m,n\leq 2\ell-1,m+n\geq 3\ell$
Table 5. Conjectured equality and failure of $b^{\ell}\left(n\right)b^{\ell}\left(m\right)\geq b^{\ell}\left(n+m\right)$ for $n\geq m\geq 1$ and $\ell\geq 5$
In the following we prove that all these cases of equality and failure in Table 5 occur. For $1\leq s\leq\ell$ and $\left(s-1\right)\ell\leq n\leq s\ell-1$ holds $b^{\ell}\left(n\right)=s$. For $0\leq r\leq\ell-1$ also $b^{\ell}\left(s\ell+r\right)=b^{\ell}\left(s\ell\right)$ holds true for all $s\geq 0$ since the smallest part in a partition different from $1$ is $\ell$. For the small values $1\leq m\leq\ell-1$ we obtain therefore $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=b^{\ell}\left(n\right)=b^{\ell}\left(n+m\right)$ if $n\equiv 0,1,\ldots,\ell-m-1\pmod{\ell}$. On the other hand $b^{\ell}\left(\ell s-1\right)<b^{\ell}\left(\ell s\right)$. Therefore, $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=b^{\ell}\left(n\right)<b^{\ell}\left(n+m\right)$ if $n\equiv\ell-m,\ell-m+1,\ldots,\ell-1\pmod{\ell}$. Furthermore for $\ell\leq n,m\leq 2\ell-1$ we obtain $2\ell\leq n+m\leq 4\ell-2$ and therefore $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=4>3=b^{\ell}\left(n+m\right)$ if $n+m\leq 3\ell-1$ and $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=4=b^{\ell}\left(n+m\right)$ if $n+m\geq 3\ell$.
Now consider generally $\left(s-1\right)\ell\leq m\leq s\ell-1$ and $\left(t-1\right)\ell\leq n\leq t\ell-1$ with $n+m\leq\ell^{2}-1$. Then $b^{\ell}\left(n\right)b^{\ell}\left(m\right)=ts$ and $b^{\ell}\left(n+m\right)=t+s$ if $n+m\geq\left(t+s-1\right)\ell$ and $b^{\ell}\left(n+m\right)=t+s-1$ if $n+m\leq\left(t+s-1\right)\ell-1$. Now we obtain $ts=t+s$ if and only if $\left(t-1\right)\left(s-1\right)=ts-t-s+1=1$ i. e. $t=2=s$ and ts=t+s1ts=t+s-1 if and only if $\left(t-1\right)\left(s-1\right)=ts-t-s+1=0$ i. e. $t=1$ or $s=1$. Similarly $ts<t+s$ if and only if $\left(t-1\right)\left(s-1\right)<1$ i. e. $t=1$ or $s=1$ and ts<t+s1ts<t+s-1 if and only if $\left(t-1\right)\left(s-1\right)<0$ i. e. never.
With PARI/GP we verified the conjecture for $\ell=6$ and $m,n\leq 3500$ and for $\ell=7$ and $m,n\leq 1200$.
Example 14.
Consider the following sequence consisting of $\alpha\left(n\right)=2^{n}+1$ for $n$ odd and $\alpha\left(n\right)=2^{n}+2$ for $n$ even. Then for $n$ odd we obtain
$\left(\alpha\left(n\right)\right)^{2}-\alpha\left(n-1\right)\alpha\left(n+1\right)$ $=$ $\left(2^{n}+1\right)^{2}-\left(2^{n-1}+2\right)\left(2^{n+1}+2\right)$
$=$ $2^{2n}+2^{n+1}+1-2^{2n}-2^{n+2}-2^{n}-4$
$=$ $-3\cdot 2^{n}-3<0$
which fulfills the log-convexity condition and for $n$ even we obtain
$\left(\alpha\left(n\right)\right)^{2}-\alpha\left(n-1\right)\alpha\left(n+1\right)$ $=$ $\left(2^{n}+2\right)^{2}-\left(2^{n-1}+1\right)\left(2^{n+1}+1\right)$
$=$ $2^{2n}+2^{n+2}+4-2^{2n}-2^{n-1}-2^{n+1}-1$
$=$ $3\cdot 2^{n-1}+3>0$
which fulfills the log-concavity condition. Therefore, the sequence is neither log-concave nor log-convex. Consider now for arbitrary $n$ $\left(\alpha\left(n\right)\right)^{n+1}\geq\left(2^{n}+1\right)^{n+1}=\left(2^{n}+1\right)^{n}\left(2^{n}+1\right)>\left(2^{n}+1\right)^{n}2^{n}=\left(2^{n+1}+2\right)^{n}\geq\left(\alpha\left(n+1\right)\right)^{n}$. For all $n$ therefore $\left(\alpha\left(n\right)\right)^{\frac{1}{n}}>\left(\alpha\left(n+1\right)\right)^{\frac{1}{n+1}}$ holds true. We can check that
$\alpha\left(n\right)\alpha\left(m\right)\geq\left(2^{n}+1\right)\left(2^{m}+1\right)=2^{n+m}+2^{n}+2^{m}+1>2^{n+m}+2\geq\alpha\left(n+m\right)$
for all $n,m$.
To summarize we have a sequence that is neither log-concave nor log-convex, satisfies the Bessenrodt–Ono inequality, and for which $\left(\alpha\left(n\right)\right)^{1/n}>\left(\alpha\left(n+1\right)\right)^{1/\left(n+1\right)}$.
Declarations and statements
Funding
This research received no external funding.
Competing interests
The authors declare that they have no competing interests.
Compliance with ethical standards
The authors declare that they comply with ethical standards
Data availability statement
No datasets were generated or analysed during the current study.
Author’s contribution statement
Both authors contributed in equal parts.